40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 4 18:13:01 2020, Yehonathan, Update, PSL, Bench testing of PSL ai channels c1psl_wire_testing_-_By_Connector.pdf
    Reply  Wed Feb 5 08:57:11 2020, Yehonathan, Update, PSL, Bench testing of PSL ai channels 
       Reply  Wed Feb 5 21:04:10 2020, Yehonathan, Update, PSL, Bench testing of PSL ai channels 
Message ID: 15189     Entry time: Wed Feb 5 21:04:10 2020     In reply to: 15187
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: Bench testing of PSL ai channels 

{Yehonathan, Jon}

We retested the failed ai channels. Most of them got fixed by applying the inverse calibration in the yaml file.

We still find some anomalous channels, mostly in the DB25 connector. Turns out, their limits were ill-defined in the EPICS database. Specifying the right limit fixed the issue.

We find one miswired channel (BNC4). We connected the BNC to the right channel on the Acromag unit which fixed the issue.

Overall all the ai channels were successfully bench-tested.


I checked the failed channels against the EPICS database definitions and the yaml file inputted to IFOTest. The channels where the readings are something other than +10/0 V, but the high/low values do change, I think can be attributed to one of two things:

  • An incorrect gain and/or offset conversion parameter in the yaml file
  • The EPICS SMOO parameter (smoothing) is set to some long value

I fixed the channel gains/offsets in the master yaml file (PSL.yaml). I also disabled smoothing in the EPICS defintions of the new PSL channels for the purpose of testing. We can uncomment those lines after installing the new chassis if noise is a problem. Please go ahead and re-test the channels again.


I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

ELOG V3.1.3-